What is the surface area to volume ratio of the rectangular prism? 4 am 10 om 8 cm 40 19 19 0 40

media%2Fe9b%2Fe9ba7078-6e08-42bf-96e2-84

media%2F4c7%2F4c771326-ec70-4170-9af6-6b
Answer #1

Please post 1 question per post, as per forum rules:

I have had opportunities to work at onsite/client locations which have helped to not only hone my technical and personal skills but bring value

Surface Area = 2(lb+bh+hl) = 2*(8*10 + 10*4 + 4*8) = 304
Volume = lbh = 4*10*8 = 320

So, Volume to Surface Area = 320/304 = 20/19